In order to find the strength of a sample of H2SO4,10g of it were taken and a piece of marble weighing 7g placed in it.When the reaction is ceased,the marble piece was removed ,dried and was found to weigh 2.2g.What is the percent strength of sulphuric acid?
 
 

Grade:11

1 Answers

Bidisha Baruah
8 years ago
mass of the marble=7g
mass of the marble left unused=2.2g
hence,mass of the marble reacted=7.0g-2.2g=4.8g
now,according to the reaction invovled in the q.100g of CaCO3 reacts with 98g of h2so4 to form caso4,h2o and CO2
so,100g of marble react with=98g of H2SO4
hence 4.8g of marble will react with=98/100x4.8=4.704g
since 10 of dil.H2so4 will contain pure h2so4=4.704g
hence 100g of dil.h2so4 will contain pure h2so4=4.704/10x100=47.04g
thus,the % strength of h2so4 is 47.04